please solve the system:
x+y=10
y=4x

Answer by Edwin McCravy(20056)

please solve the system:

  x + y = 10
      y = 4x

The second equation is already solved for y so
substitute 4x in place of y in the first:

  x + y = 10

Putting 4x where y is:

 x + 4x = 10
1x + 4x = 10
     5x = 10
      x = 2

Now substitute 2 for x in the equation y = 4x

      y = 4x
      y = 4(2)
      y = 8

So the solution is (x,y) = (2,8)

Edwin